# Provable Benefit of Annealed Langevin Monte Carlo for Non-log-concave Sampling

### v1

@@ 5 Analysis of Annealed Langevin Monte Carlo
It is crucial to recognize that the ALD (Equation 4) cannot be precisely simulated in practice. Transitioning from LD to LMC introduces additional errors due to discretization. This section will present a detailed convergence analysis for the annealed Langevin Monte Carlo (ALMC) algorithm, which is implementable in practical scenarios.
A straightforward yet non-optimal method to discretize Equation 4 involves employing the Euler-Maruyama scheme, i.e.,
@@
However, considering that $\nabla\log\widetilde{\pi}_{t}(x)=-\eta\left(\frac{t}{T}\right)\nabla V(x)-\lambda\left(\frac{t}{T}\right)x$, the integral of the linear term can be computed in closed form, so we can use the exponential-integrator scheme [zhang2023fast, zhang2023gddim] to further reduce the discretization error.
Given the total time $T$, we define a sequence of points $10=\theta_{0}<\theta_{1}<...<\theta_{M}=$ 1, and set $T_{\ell}=T\theta_{\ell}$, $h_{\ell}=T(\theta_{\ell}-\theta_{\ell-1})$. The exponential-integrator scheme is then expressed as
${\rm d}X_{t}=\left(-\eta\left(\frac{t}{T}\right)\nabla V(X_{t_{-}})-\lambda\left(\frac{t}{T}\right)X_{t}\right){\rm d}t+\sqrt{2}{\rm d}B_{t},~{}t\in[0,T],~{}X_{0}\sim\pi_{0},$ (6)
where $t_{-}:=T_{\ell-1}$ when $t\in[T_{\ell-1},T_{\ell})$, $\ell\in\left[\!\left[1,M\right]\!\right]$. The explicit update rule is detailed in Algorithm 1, with $x_{\ell}$ denoting $X_{T_{\ell}}$, and the derivation of Equation 7 is presented in Section B.1. Notably, replacing $\nabla V(X_{t_{-}})$ with $\nabla V(X_{t})$ recovers the ALD (Equation 4), and setting $\eta\equiv 1$ and $\lambda\equiv 0$ reduces to the classical LMC iterations.
@@
1 For $0\leq\theta<\theta^{\prime}\leq 1$, define $\Lambda_{0}(\theta^{\prime},\theta)=\exp\left(-T\int_{\theta}^{\theta^{\prime}}\lambda(u){\rm d}u\right)$, $H(\theta^{\prime},\theta)=T\int_{\theta}^{\theta^{\prime}}\eta(u)\Lambda_{0}(u,\theta^{\prime}){\rm d}u$, and $\Lambda_{1}(\theta^{\prime},\theta)=\sqrt{2T\int_{\theta}^{\theta^{\prime}}\Lambda_{0}^{2}(u,\theta^{\prime}){\rm d}u}$;
@@
3Obtain a sample $x_{0}\sim\pi_{0}$ using rejection sampling (Algorithm 2);
@@ 7       Update
We illustrate the ALMC algorithm in Figure 1. The underlying intuition behind this algorithm is that by setting a sufficiently long total time $T$, the trajectory of the continuous dynamic (i.e., annealed LD) approaches the reference trajectory closely, as established in Theorem 1. Additionally, by adopting sufficiently small step sizes $h_{1},...,h_{M}$, the discretization error can be substantially reduced. Unlike traditional annealed LMC methods, which require multiple LMC update steps for each intermediate distribution $\pi_{1},...,\pi_{M}$, our approach views the annealed LMC as a discretization of a continuous-time process. Consequently, it is adequate to perform a single step of LMC towards each intermediate distribution $\widetilde{\pi}_{T_{1}},...,\widetilde{\pi}_{T_{M}}$, provided that $T$ is sufficiently large and the step sizes are appropriately small.
@@
Figure 1: Illustration of the ALMC algorithm. The $\ell$-th green arrow, proceeding from left to right, represents one step of LMC towards $\widetilde{\pi}_{T_{\ell}}$ with step size $h_{\ell}$, while each red arrow corresponds to the application of the same transition kernel, initialized at $\widetilde{\pi}_{T_{\ell-1}}$ on the reference trajectory $\mathbb{P}$, which is depicted in purple. To evaluate $\operatorname{KL}\left(\mathbb{P}\middle\|\mathbb{Q}\right)$, the Girsanov theorem implies that we only need to bound the aggregate KL divergence across each small interval (i.e., the sum of the blue “distances”).
The subsequent theorem provides a convergence guarantee for the annealed LMC algorithm, with the complete proof detailed in Section B.2.
@@
Under Assumptions 1 and 2, Algorithm 1 can generate a distribution $\nu^{\rm ALMC}$ satisfying $\operatorname{KL}\left(\pi\middle\|\nu^{\rm ALMC}\right)\leq\varepsilon^{2}$ within
$\widetilde{O}\left(\frac{d\beta^{2}{\cal A}_{\eta,\lambda}(\pi)^{2}}{\varepsilon^{6}}\right)$
@@
Let $\mathbb{Q}$ be the path measure of ALMC (Equation 4), whose marginal distribution at time $T$ is the output distribution $\nu^{\rm ALMC}$. Again, let $\mathbb{P}$ denote the reference path measure of Equation 5 used in the proof of Theorem 1, in which the same vector field $(v_{t})_{t\in[0,T]}$ ensures that $X_{t}\sim\widetilde{\pi}_{t}$ under $\mathbb{P}$ for all $t\in[0,T]$.
Applying Girsanov theorem (Lemma 1) and carefully dealing with the discretization error, we can upper bound $\operatorname{KL}\left(\mathbb{P}\middle\|\mathbb{Q}\right)$ by
@@
$\operatorname{KL}\left(\mathbb{P}\middle\|\mathbb{Q}\right)\lesssim\sum_{\ell=1}^{M}\left(\frac{1+\eta(\theta_{\ell})^{2}\beta^{2}h_{\ell}^{2}}{T}\int_{\theta_{\ell-1}}^{\theta_{\ell}}\left|\dot{\pi}\right|^{2}_{\theta}{\rm d}\theta+\eta(\theta_{\ell})^{2}\beta^{2}dh_{\ell}^{2}\left(1+h_{\ell}\left(\beta\eta(\theta_{\ell})+\lambda(\theta_{\ell-1})\right)\right)\right).$
@@
The first summation is governed by the total time $T$, which pertains to the convergence of the continuous dynamic (i.e., ALD). Setting $T\asymp\frac{{\cal A}_{\eta,\lambda}(\pi)}{\varepsilon^{2}}$ ensures that the first summation remains $O(\varepsilon^{2})$, provided that the step size $h_{\ell}$ is sufficiently small. The second summation addresses the discretization error, and it suffices to determine the appropriate value of $\theta_{\ell}$ to minimize $M$, the total number of calls to the oracle of $\nabla V$ for discretizing the SDE. Combining $M$ with the complexity of sampling from $\pi_{0}$ determines the overall complexity of the algorithm. ∎
Once again, our analysis relies on bounding the global error between two path measures by Girsanov theorem. The annealing schedule $\eta(\cdot)$ and $\lambda(\cdot)$ influence the complexity exclusively through the action ${\cal A}_{\eta,\lambda}(\pi)$. Identifying the optimal annealing schedule to minimize this action remains a significant challenge, and we propose this as an area for future research.
@@
We also note that our Assumption 1 encompasses strongly-log-concave target distributions. For sampling from these well-conditioned distributions via LMC, the complexity required to achieve $\varepsilon$- in TV distance scales as $\widetilde{O}\left(\varepsilon^{-2}\right)$ [chewi2022log]. However, using Pinsker inequality ${\rm KL}\geq 2{\rm TV}^{2}$, our complexity to meet the same error criterion is $O\left(\varepsilon^{-6}\right)$, indicating a significantly higher computational demand. Refining the parameter dependence in our algorithm to reduce this complexity remains a key area for future work.
Finally, we present a conjecture concerning the action ${\cal A}_{\eta,\lambda}(\pi)$, leaving its proof or disproof as an open question for future research. Following this conjecture, we demonstrate its applicability to a specific class of non-log-concave distributions in the subsequent example, with the detailed proof provided in Appendix C.
Conjecture 1.
Under Assumption 1, the action ${\cal A}_{\eta,\lambda}(\pi)$ is bounded above by a polynomial function of the problem parameters, including $d$, $\beta$, $R$, $\operatorname{\mathbb{E}}_{\pi}\left[\left\|\cdot\right\|^{2}\right]$ $\eta_{0}$, $\lambda_{0}$, etc.
@@
Consider a mixture of Gaussian distributions in $\mathbb{R}^{d}$ defined by $\pi=\sum_{i=1}^{N}p_{i}\operatorname{{\cal N}}\left(y_{i},\beta^{-1}I\right)$, where the weights $p_{i}>0$, $\sum_{i=1}^{N}p_{i}=1$, and $\left\|y_{i}\right\|=r$ for all $i\in\left[\!\left[1,N\right]\!\right]$. Consequently, the potential $V=-\log\pi$ is $B$-smooth, where $B=\beta(4r^{2}\beta+1)$. With an annealing schedule defined by $\eta(\cdot)\equiv 1$ and $\lambda(\theta)=dB(1-\theta)^{\gamma}$ for some $1\leq\gamma=O(1)$, it follows that ${\cal A}_{\eta,\lambda}(\pi)=O\left(d(r^{2}\beta+1)\left(r^{2}+\frac{d}{\beta}\right)\right)$.
To demonstrate the superiority of our theoretical results, let us consider a simplified scenario where $N=2$, $y_{1}=-y_{2}$, and $r^{2}\gg\frac{1}{\beta}$. We derive from Example 1 and Pinsker inequality that the total complexity required to obtain a sample that is $\varepsilon$-accurate in TV is $\widetilde{O}\left(d^{3}\beta^{2}r^{4}(r^{4}\beta^{2}\vee d^{2})\varepsilon^{-6}\right)$. In contrast, studies such as [schlichting2019poincare, chen2021dimensionfree, dong2022spectral] indicate that the LSI constant of $\pi$ is $\Omega\left({\rm e}^{\Theta(\beta r^{2})}\right)$, so existing bounds in Table 1 suggest that LMC, to achieve the same accuracy, can only provide an exponential complexity guarantee of $\widetilde{O}\left({\rm e}^{\Theta(\beta r^{2})}d\varepsilon^{-2}\right)$.

### v2

@@ 5 Analysis of Annealed Langevin Monte Carlo
It is crucial to recognize that, in practice, running the algorithm requires knowledge of the score functions $(\nabla\log\pi_{\theta})_{\theta\in[0,1]}$ in closed form. Additionally, simulating ALD (Eq. 3) necessitates discretization schemes, which introduces further errors. This section presents a detailed non-asymptotic convergence analysis for the annealed Langevin Monte Carlo (ALMC) algorithm, which is a practical approach for real-world implementations.
The outline of this section is as follows. We will first introduce a family of the interpolation curve $(\pi_{\theta})_{\theta\in[0,1]}$ that serves as the basis for our analysis. Next, we propose a discretization scheme tailored to the special structure of $(\pi_{\theta})_{\theta\in[0,1]}$. Finally, we state the complexity analysis in Thm. 2, and provide an example demonstrating the improvement of our bounds over existing results.
5.1 Choice of the Interpolation Curve
We consider the following curve of probability measures on $\mathbb{R}^{d}$:
$\pi_{\theta}\propto\exp\left(-\eta(\theta)V-\frac{\lambda(\theta)}{2}\|\cdot\|^{2}\right),~{}\theta\in[0,1],$ (5)
where the functions $\eta(\cdot)$ and $\lambda(\cdot)$ are called the annealing schedule. These functions must be differentiable and monotonic, satisfying the boundary conditions $\eta_{0}=\eta(0)\nearrow\eta(1)=1$ and $\lambda_{0}=\lambda(0)\searrow\lambda(1)=0$. The values of $\eta_{0}\in[0,1]$ and $\lambda_{0}\in(1,{+\infty})$ are chosen so that $\pi_{0}$ corresponds a Gaussian distribution $\operatorname{{\cal N}}\left(0,\lambda_{0}^{-1}I\right)$ or can be efficiently sampled within $\widetilde{O}(1)$ steps of rejection sampling (see Lem. 5 for details). We also remark that, under Assump. 2, $\pi_{\theta}$ has a finite second-order moment (see Lem. 8 the proof), ensuring the $\text{W}_{2}$ distance between $\pi_{\theta}$’s is well-defined.
This flexible interpolation scheme includes many general cases. For example, Brosse et al. (2018) and Ge et al. (2020) used the schedule $\eta(\cdot)\equiv 1$, while Neal (2001) used the schedule $\lambda(\theta)=c(1-\eta(\theta))$. The key motivation for this interpolation is that when $\theta=0$, the quadratic term predominates, making the potential of $\pi_{0}$ strongly-convex with a moderate condition number, thus $\pi_{0}$ is easy to sample from; on the other hand, when $\theta=1$, $\pi_{1}$ is just the target distribution $\pi$. As $\theta$ gradually increases from 00 to $1$, the readily sampleable distribution $\pi_{0}$ slowly transforms into the target distribution $\pi_{1}$.
5.2 The Discretization Algorithm
With the interpolation curve $(\pi_{\theta})_{\theta\in[0,1]}$ chosen as above, a straightforward yet non-optimal method to discretize Eq. 3 involves employing the Euler-Maruyama scheme, i.e.,
@@ 5.2 The Discretization Algorithm
However, considering that $\nabla\log\widetilde{\pi}_{t}(x)=-\eta\left(\frac{t}{T}\right)\nabla V(x)-\lambda\left(\frac{t}{T}\right)x$, the integral of the linear term can be computed in closed form, so we can use the exponential-integrator scheme (Zhang & Chen, 2023; Zhang et al., 2023) to further reduce the discretization error. Given the total time $T$, we define a sequence of points $10=\theta_{0}<\theta_{1}<...<\theta_{M}=$ 1, and set $T_{\ell}=T\theta_{\ell}$, $h_{\ell}=T(\theta_{\ell}-\theta_{\ell-1})$. The exponential-integrator scheme is then expressed as
$\mathrm{d}X_{t}=\left(-\eta\left(\frac{t}{T}\right)\nabla V(X_{t_{-}})-\lambda\left(\frac{t}{T}\right)X_{t}\right)\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},~{}t\in[0,T],~{}X_{0}\sim\pi_{0},$ (6)
where $t_{-}:=T_{\ell-1}$ when $t\in[T_{\ell-1},T_{\ell})$, $\ell\in\left[\!\left[1,M\right]\!\right]$. The explicit update rule is detailed in Alg. 1, with $x_{\ell}$ denoting $X_{T_{\ell}}$, and the derivation of Eq. 7 is presented in Sec. C.1. Notably, replacing $\nabla V(X_{t_{-}})$ with $\nabla V(X_{t})$ recovers the ALD (Eq. 3), and setting $\eta\equiv 1$ and $\lambda\equiv 0$ reduces to the classical LMC iterations.
@@
1 For $0\leq\theta<\theta^{\prime}\leq 1$, define $\Lambda_{0}(\theta^{\prime},\theta)=\exp\left(-T\int_{\theta}^{\theta^{\prime}}\lambda(u)\mathrm{d}u\right)$, $H(\theta^{\prime},\theta)=T\int_{\theta}^{\theta^{\prime}}\eta(u)\Lambda_{0}(u,\theta^{\prime})\mathrm{d}u$, and $\Lambda_{1}(\theta^{\prime},\theta)=\sqrt{2T\int_{\theta}^{\theta^{\prime}}\Lambda_{0}^{2}(u,\theta^{\prime})\mathrm{d}u}$;
@@
3Obtain a sample $x_{0}\sim\pi_{0}$ (e.g., using rejection sampling (Alg. 2));
@@ 7       Update
We illustrate the ALMC algorithm in Fig. 1. The underlying intuition behind this algorithm is that by setting a sufficiently long total time $T$, the trajectory of the continuous dynamic (i.e., annealed LD) approaches the reference trajectory closely, as established in Thm. 1. Additionally, by adopting sufficiently small step sizes $h_{1},...,h_{M}$, the discretization error can be substantially reduced. Unlike traditional annealed LMC methods, which require multiple LMC update steps for each intermediate distribution $\pi_{1},...,\pi_{M}$, our approach views the annealed LMC as a discretization of a continuous-time process. Consequently, it is adequate to perform a single step of LMC towards each intermediate distribution $\widetilde{\pi}_{T_{1}},...,\widetilde{\pi}_{T_{M}}$, provided that $T$ is sufficiently large and the step sizes are appropriately small.
@@
Figure 1: Illustration of the ALMC algorithm. The $\ell$-th green arrow, proceeding from left to right, represents one step of LMC towards $\widetilde{\pi}_{T_{\ell}}$ with step size $h_{\ell}$, while each red arrow corresponds to the application of the same transition kernel, initialized at $\widetilde{\pi}_{T_{\ell-1}}$ on the reference trajectory $\mathbb{P}$, which is depicted in purple. To evaluate $\operatorname{KL}(\mathbb{P}\|\mathbb{Q})$, the Girsanov theorem implies that we only need to bound the aggregate KL divergence across each small interval (i.e., the sum of the blue “distances”).
5.3 Convergence Analysis of the Algorithm
The subsequent theorem provides a convergence guarantee for the annealed LMC algorithm, with a detailed proof available in Sec. C.2.
@@ 5.3 Convergence Analysis of the Algorithm
Under Assumps. 2 and 1, Alg. 1 can generate a distribution $\nu^{\rm ALMC}$ satisfying $\operatorname{KL}(\pi\|\nu^{\rm ALMC})\leq\varepsilon^{2}$ within
$\widetilde{O}\left(\frac{d\beta^{2}{\cal A}^{2}}{\varepsilon^{6}}\right)$
@@
Let $\mathbb{Q}$ be the path measure of ALMC (Eq. 3), the time-discretized sampling process, whose marginal distribution at time $T$ is the output distribution $\nu^{\rm ALMC}$. Again, let $\mathbb{P}$ denote the reference path measure of Eq. 4 used in the proof of Thm. 1, in which the same vector field $(v_{t})_{t\in[0,T]}$ ensures that $X_{t}\sim\widetilde{\pi}_{t}$ under $\mathbb{P}$ for all $t\in[0,T]$. Applying Girsanov theorem (Lem. 1) and carefully dealing with the discretization error, we can upper bound $\operatorname{KL}(\mathbb{P}\|\mathbb{Q})$ by
@@
$\operatorname{KL}(\mathbb{P}\|\mathbb{Q})\lesssim\sum_{\ell=1}^{M}\left(\frac{1+\eta(\theta_{\ell})^{2}\beta^{2}h_{\ell}^{2}}{T}\int_{\theta_{\ell-1}}^{\theta_{\ell}}|\dot{\pi}|^{2}_{\theta}\mathrm{d}\theta+\eta(\theta_{\ell})^{2}\beta^{2}dh_{\ell}^{2}\left(1+h_{\ell}\left(\beta\eta(\theta_{\ell})+\lambda(\theta_{\ell-1})\right)\right)\right).$
@@
The first summation is governed by the total time $T$, which pertains to the convergence of the continuous dynamic (i.e., ALD). Setting $T\asymp\frac{{\cal A}}{\varepsilon^{2}}$ ensures that the first summation remains $O(\varepsilon^{2})$, provided that the step size $h_{\ell}$ is sufficiently small. The second summation addresses the discretization error, and it suffices to determine the appropriate value of $\theta_{\ell}$ to minimize $M$, the total number of calls to the oracle of $\nabla V$ for discretizing the SDE. Combining $M$ with the complexity of sampling from $\pi_{0}$ determines the overall complexity of the algorithm. ∎
Once again, our analysis relies on bounding the global error between two path measures by Girsanov theorem. The annealing schedule $\eta(\cdot)$ and $\lambda(\cdot)$ influences the complexity exclusively through the action ${\cal A}$. This crucial identity, based on the choice of interpolation curve $(\pi_{\theta})_{\theta\in[0,1]}$, significantly affects the effectiveness of ALMC. Our specific choice (Eq. 5), among all interpolation curves with closed-form scores, aims at simplifying the discretization error analysis. However, alternative annealing schemes could be explored to find the optimal one in practice. For instance, a recent paper (Fan et al., 2024) proposed an annealing scheme $\pi_{\theta}(x)\propto\pi_{0}((1-a\theta)x)^{1-\theta}\pi_{1}\left(\frac{x}{b+(1-b)\theta}\right)^{\theta}$ for some $a\in[0,1]$ and $b\in(0,1]$. The extra scaling factor $\frac{1}{b+(1-b)\theta}$ might improve mixing speed. We leave the choice of the interpolation curve as a topic for future research.
@@
We also note that our Assump. 2 encompasses strongly-log-concave target distributions. For sampling from these well-conditioned distributions via LMC, the complexity required to achieve $\varepsilon$-accuracy in TV distance scales as $\widetilde{O}(\varepsilon^{-2})$ (Chewi, 2024). However, using Pinsker inequality $\operatorname{KL}\geq 2\operatorname{TV}^{2}$, our complexity to meet the same error criterion is $O(\varepsilon^{-6})$, indicating a significantly higher computational demand. While our discretization error is as sharp as existing works, the main reason for our worse $\varepsilon$-dependence is due to the time required for the continuous dynamics to converge, which is based on Girsanov theorem rather than LSI. While sharpening the $\varepsilon$-dependence based on Girsanov theorem remains a challenge, our $O(\varepsilon^{-6})$ dependence still outperforms all the other bounds in Tab. 1 without isoperimetry assumptions.
Finally, we conclude by demonstrating an example of a class of mixture of Gaussian distributions, which illustrates how our analysis can improve the complexity bound from exponential to polynomial. The detailed proof is provided in App. D.
@@
Consider a mixture of Gaussian distributions in $\mathbb{R}^{d}$ defined by $\pi=\sum_{i=1}^{N}p_{i}\operatorname{{\cal N}}\left(y_{i},\beta^{-1}I\right)$, where the weights $p_{i}>0$, $\sum_{i=1}^{N}p_{i}=1$, and $\|y_{i}\|=r$ for all $i\in\left[\!\left[1,N\right]\!\right]$. Consequently, the potential $V=-\log\pi$ is $B$-smooth, where $B=\beta(4r^{2}\beta+1)$. With an annealing schedule defined by $\eta(\cdot)\equiv 1$ and $\lambda(\theta)=dB(1-\theta)^{\gamma}$ for some $1\leq\gamma=O(1)$, it follows that ${\cal A}=O\left(d(r^{2}\beta+1)\left(r^{2}+\frac{d}{\beta}\right)\right)$.
To show the superiority of our theoretical results, consider a simplified scenario with $N=2$, $y_{1}=-y_{2}$, and $r^{2}\gg\beta^{-1}$. By applying Ex. 2 and the Pinsker inequality, the total complexity required to obtain an $\varepsilon$-accurate sample in TV distance is $\widetilde{O}(d^{3}\beta^{2}r^{4}(r^{4}\beta^{2}\vee d^{2})\varepsilon^{-6})$. In contrast, studies such as Schlichting (2019); Chen et al. (2021); Dong & Tong (2022) indicate that the LSI constant of $\pi$ is $\Omega({\rm e}^{\Theta(\beta r^{2})})$, so existing bounds in Tab. 1 suggest that LMC, to achieve the same accuracy, can only provide an exponential complexity guarantee of $\widetilde{O}({\rm e}^{\Theta(\beta r^{2})}d\varepsilon^{-2})$.
6 Experiments
We conduct simple numerical experiments to verify the findings in Ex. 2, which demonstrate that for a certain class of mixture of Gaussian distributions, ALMC achieves polynomial convergence with respect to $r$. Specifically, we consider a target distribution comprising a mixture of $6$ Gaussians with uniform weights $\frac{1}{6}$, means $\left\{\left(r\cos\frac{k\pi}{3},r\sin\frac{k\pi}{3}\right):k\in\left[\!\left[0,5\right]\!\right]\right\}$, and the same covariance $0.1I$ (corresponding to $\beta=10$ in Ex. 2). We experimented $r\in\{2,5,10,15,20,25,30\}$, and compute the number of iterations required for the empirical KL divergence from the target distribution to the sampled distribution to fall below $0.2$ (blue curve) and $0.1$ (orange curve). The results, displayed in Fig. 2, use a log-log scale for both axes. The near-linear behavior of the curves in this plot confirms that the complexity depends polynomially on $r$, validating our theory. Further details of the experimental setup and implementation can be found in App. F.
Figure 2: Relationship between norm of means $r$ and the number of iterations to reach $0.2$ and $0.1$ accuracy in KL divergence, both in $\log$ scale.